# Quantum Localization in Incommensurate Tight-Binding Chains

### v1

@@ I Introduction
Electronic properties of disordered and quasi-periodic materials have long been the focus of extensive research in condensed matter physics. In periodic crystals, where atoms are arranged in regular, repeating patterns, the wave function can be expressed as the product of a plane wave and a function with the same periodicity as the lattice, a result known as Bloch’s theorem Bloch (1929). Physically, this implies that the probability distribution of the electron is spread evenly throughout the lattice. However, in disordered or quasi-periodic materials, this result does not hold. In such systems, it is possible for the wave functions of electrons to become localized.
@@
One of the first theoretical frameworks developed to study disordered materials is the Anderson model Anderson (1958); Lee and Ramakrishnan (1985); Kramer and MacKinnon (1993a). In this model, the transfer of electrons between nearest-neighbor sites of a crystal lattice is described using the tight-binding approximation Ashcroft and Mermin (1976); Kittel (2004), while the on-site energies are randomly selected from a specified range, with the extent of this range referred to as the disorder strength. In one dimension (1D), all electronic states become exponentially localized regardless of the disorder strength Mott and Twose (1961). Various methods have been employed to analyze the Anderson model, including the transfer matrix methods Mackinnon and Kramer (1983), Green’s function techniques Thouless (1972), and numerical simulations such as exact diagonalization and the Lanczos algorithm Kramer and MacKinnon (1993b). These techniques calculate key physical quantities such as the localization length, density of states, and conductivity. Experimental validations of the Anderson model have been conducted using ultra-cold atom systems Billy et al. (2008) and photonic lattices Schwartz et al. (2007).
@@
Localization in disordered systems usually tends to be suppressed by an external magnetic field. For instance, weakly localized systems in two dimensions (2D) exhibit negative magnetoresistance due to a reduction of constructive interference of the electronic wave functions caused by an additional magnetic phase shift Bergmann (1984). In the 2D Anderson model, a weak magnetic field reduces localization for states near the Landau band centers Ando (1989).
@@
A different class of systems, which are not randomly disordered but still lack perfect lattice periodicity, are quasi-periodic crystals. One of the most popular quasi-periodic 1D models is the Aubry-André (AA) model Aubry and André (1980); Sokoloff (1985); Domínguez-Castro and Paredes (2019), which features periodically varying on-site energies with the period incommensurate with the lattice, meaning the ratio between the two periods is irrational. In the AA model, all states become localized when the potential strength exceeds a certain threshold. The model is closely related to the Harper model, which describes electrons in a two-dimensional periodic lattice under a uniform magnetic field; the Harper equation can be mapped onto the AA model Harper (1955). The AA model can also be used to describe ultracold atoms in two incommensurate optical lattices. An experimental demonstration of the localization of bosonic atoms in such a system was reported in Ref. Roati et al. (2008), offering direct evidence of the AA model predictions. Further experiments have investigated many-body localization in quasi-periodic systems Schreiber et al. (2015).
@@
Several models related to the AA model have been proposed to explore more complex localization phenomena in quasi-periodic systems. One such model is the Soukoulis–Economou model Soukoulis and Economou (1982), which introduces an additional potential term that breaks the self-duality of the original AA model. This modification allows the coexistence of both extended and localized states within the same system, separated by a mobility edge—critical energy value marking the transition between extended and localized behavior. Mobility edges have also been found theoretically in numerous other generalizations of the single-chain AA model Das Sarma et al. (1988); Biddle et al. (2009); Biddle and Das Sarma (2010); Liu et al. (2015); Cestari et al. (2016); Liu and Guo (2018); Wang et al. (2020); Vu and Das Sarma (2023). Experimental evidence of a mobility edge in a 1D quasiperiodic optical lattice has been reported in Ref. Lüschen et al. (2018).
@@
The Fibonacci chain Kohmoto and Banavar (1986); Fujita and Machida (1987); Kohmoto et al. (1987); Jagannathan (2021) provides yet another important example. In this 1D tight-binding model, the on-site potentials or hopping amplitudes follow a deterministic, quasi-periodic pattern based on the Fibonacci sequence. The resulting wavefunctions are critical, i.e. neither fully extended nor exponentially localized, instead decaying algebraically and exhibiting self-similar, fractal features. Interestingly, it has been shown that the AA model and the Fibonacci quasi-crystal are topologically equivalent, as one can be smoothly deformed into the other without closing any bulk energy gaps Kraus and Zilberberg (2012).
@@
In this paper, we construct a model of quasi-periodic crystals in which (i) incommensurability naturally arises from the geometry of the system and (ii) it is straightforward to study the effects of an external magnetic field on quasi-1D localization. Both goals can be achieved if one considers two coupled 1D chains of atoms, with the ratio $\rho$ of the lattice periods being an irrational number. We do not introduce any external on-site potentials, therefore, in contrast to the AA model and its multi-chain generalizations Guo et al. (2014); Rossignolo and Dell’Anna (2019); Wang et al. (2021); Lin et al. (2023); Aghtouman and Hosseini (2024), the incommensurability in our model affects only the hopping amplitudes.
@@
For numerical analysis, we assume that the chains have different numbers of atoms but the same length, with periodic boundary conditions, and approximate $\rho$ by a ratio of two large coprime numbers. To characterize localization of the electronic wave functions, we use the the inverse participation ratio (IPR) Bell and Dean (1970); Thouless (1974); Wegner (1980). The IPR takes values between zero and one, with the larger values corresponding to the more localized wave functions.
@@ II The Model
to closely approximate an irrational number. This can be achieved using the method of continued fractions Azbel (1979); Sokoloff (1981). We represent the ratio (1) in the following form:
$\rho=a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{\ddots+\cfrac{1}{a_{p}}}},$ (2)
where $a_{0},a_{1},\ldots,a_{p}$ are integers Khinchin (1964). In our analysis we will try to bring $\rho$ close to the golden ratio $\phi=(1+\sqrt{5})/2\simeq 1.618$, which corresponds to all $a_{i}=1$ ($i\geq 0$). Truncating this infinite sequence to $p$ terms leads to a rational approximation for $\phi$ called the $p$th convergent and denoted as $\phi_{p}$. The latter is given by a ratio of consecutive Fibonacci numbers:
$\phi_{p}=\frac{F_{p+1}}{F_{p}}.$ (3)
@@
see Fig. 1b. Therefore, the hopping amplitudes take the following form:
@@
Note that $t_{A}$ and $t_{B}$ are invariant with respect to the lattice translations in the respective chains, i.e., depend only on $|i_{A}-j_{A}|$ or $|i_{B}-j_{B}|$, but have different magnitudes if $N_{A}\neq N_{B}$. In contrast, the inter-chain hopping amplitudes are not translationally invariant.
Although all possible hopping terms are included in our model, in practice, for any given atom, only a few neighbors in both chains make a non-negligible contribution. The magnitudes of the inter-chain hoppings vary from atom to atom in a quasi-random fashion, producing a “geometrical disorder” in the system. We would like to mention here the study of the energy spectrum in a model of two coupled weakly incommensurate “Moiré chains” Carr et al. (2020). In this model, only the nearest-neighbor intra-chain hopping amplitudes (same in both chains) were considered and a different dependence of the inter-chain hoppings on the distance between the sites was assumed.
@@ II.1 Effects of magnetic field
We now include the effects of an external magnetic field in our tight-binding model, which can be achieved by using the Peierls substitution Hofstadter (1976); Marder (2010). In this method each hopping amplitude in Eqs. (5) and (6) is modified by a complex phase factor: $t_{\bm{r}\bm{r}^{\prime}}\xrightarrow{}t_{\bm{r}\bm{r}^{\prime}}e^{i\varphi_{\bm{r}\bm{r}^{\prime}}}$, where
@@ III Numerical Analysis
In periodic systems, Bloch’s theorem enables analytical calculation of the energy spectrum by assuming extended plane wave solutions for the wave function. The number of atoms per unit cell in a periodic lattice directly corresponds to the number of energy bands Simon (2013). However, our system does not repeat periodically; instead, it consists of $N_{A}+N_{B}$ atoms within a single “unit cell”, which coincides with the whole crystal. This non-periodicity means that extended plane-wave solutions are no longer valid and that Bloch’s theorem cannot be applied, resulting in $N_{A}+N_{B}$ distinct energy levels rather than energy bands. By increasing the size of the system and tuning the ratio $\rho$ to approximate an irrational number, we effectively render the system incommensurate. Since there is no analytical solution available for such a system, we must resort to numerical methods to find the wave functions and energy eigenvalues.
@@
In our model, following a considerable precedent in the previous studies of incommensurate 1D systems Aubry and André (1980), we focus on using the golden ratio $\phi$ as the incommensurability parameter. To better represent real physical systems we select large values for $N_{A}$ and $N_{B}$. By choosing values of $N_{A}$ and $N_{B}$ in the thousands, we ensure a high degree of precision in approximating $\phi$ while maintaining practical relevance for numerical analysis. Namely, we use
@@
as determined from Eq. (3), i.e. set $N_{A}=F_{18}=2584$ and $N_{B}=F_{17}=1597$, where $F_{p}$ is the $p$th Fibonacci number. Selecting higher-order approximations does not qualitatively affect our results.
@@ IV Results
In our numerical calculations, we use $N_{A}=2584$ and $N_{B}=1597$ and diagonalize a $4181\times 4181$ Hamiltonian matrix for different values of the dimensionless lattice constant $a$ and the inter-chain separation $d$ (the lattice constant in the $B$ chain is given by $b=\rho a$). The effects of a magnetic field, which are described by the dimensionless parameters $\chi_{\parallel}$ and $\chi_{\perp}$, are considered in Sec. IV.1.
@@
The evolution of the energy spectrum as a function of the inter-chain separation $d$ is shown in Fig. 5. At sufficiently large values of $d$ the spectrum is smooth and agrees with analytical results for decoupled chains. When $d$ is decreased, gaps begin to appear in the spectrum. Interestingly, these gaps seem to be located exactly at indices corresponding to the Fibonacci numbers $n=F_{p}$ or their simple combinations. This can be understood as an effect of “small denominators” in the perturbation expansions with respect to a weak inter-chain coupling, see Appendix A.
@@
Our results for the coupled chains indicate that higher-energy states tend to be localized, exhibiting nonzero values of the IPR, whereas lower-energy states remain extended. Notably, the transition from extended to localized states is not continuous; instead, it is abrupt and coincides with a gap in the energy spectrum. These findings suggest a strong correlation between the presence of energy gaps and the onset of localization in higher-energy states.
@@
To illustrate the shape of the localized states, in Fig. 7 we plot the probability densities $|\psi_{n}(A,i_{A})|^{2}$ and $|\psi_{n}(B,i_{B})|^{2}$ using as an example the state with $n=2750$. The wave function exhibits sharply localized peaks in both chains in close proximity of each other. To characterize its spatial extent, we approximate the probability density in each chain by an exponential function of the form
$|\psi(\theta)|^{2}=|\psi(\theta_{0})|^{2}\exp\left({-\frac{|\theta-\theta_{0}|}{\delta}}\right),$ (19)
where we use the angular coordinate $\theta$ to take advantage of the cylindrical symmetry of the system. By fitting the numerical data, we can extract the angular localization length $\delta$. In Fig. 7b, we show the results of this procedure for the probability density in the AA and $B$ chains separately, thereby determining two distinct localization lengths, $\delta_{A}$ and $\delta_{B}$. Note that we are not making the claim that the wave functions are exponentially localized based on our data. We use the simple fitting function (19) only to get some qualitative insight into the spatial extent of the localized eigenstates.
It may be more convenient to define a linear localization length $\xi=R\delta$, doing this we can compare the extent of localization to the lattice constant. For the $n=2750$ state, we find $\xi_{A}\simeq 0.387a$ and $\xi_{B}\simeq 0.309b$ ($0.499a$). Thus the wave function is strongly localized on the scale of the lattice constant.
Figure 7: (Color online) (a) “Bird’s-eye” view of the probability densities $|\psi_{n}(A,i_{A})|^{2}$ and $|\psi_{n}(B,i_{B})|^{2}$ in the localized state with $n=2750$, $a=2$, and $d=1$. (b) Close‐up view of the same probability densities, using cylindrical coordinates $\theta$ and $z$. Data is approximately fit to Eq. (19).
@@
Many studies have examined the behaviour of wave functions at the transition between localized and extended phases. In the Aubry-André model, for example, wave functions at the critical point are neither fully localized nor extended, but exhibit irregular multifractural fluctuations across all length scales Siebesma and Pietronero (1987). In contrast, our model shows an abrupt transition between localized and extended states, with no intermediate phase. This means that on either side of the mobility edge, the wave functions are either localized or extended, sharing the same properties as other states in their respective phases.
IV.1 Effects of magnetic field

### v2

@@ I Introduction
Electronic properties of disordered and quasi-periodic materials have long been the focus of extensive research in condensed matter physics. In periodic crystals, where atoms are arranged in regular, repeating patterns, the wave function can be expressed as the product of a plane wave and a function with the same periodicity as the lattice, a result known as Bloch’s theorem [17]. Physically, this implies that the probability distribution of the electron is spread evenly throughout the lattice. However, in disordered or quasi-periodic materials, this result does not hold. In such systems, it is possible for the wave functions of electrons to become localized.
@@
One of the first theoretical frameworks developed to study disordered materials is the Anderson model [7, 39, 36]. In this model, the transfer of electrons between nearest-neighbor sites of a crystal lattice is described using the tight-binding approximation [9, 32], while the on-site energies are randomly selected from a specified range, with the extent of this range referred to as the disorder strength. In one dimension (1D), all electronic states become exponentially localized regardless of the disorder strength [48]. Various methods have been employed to analyze the Anderson model, including the transfer matrix methods [46], Green’s function techniques [59], and numerical simulations such as exact diagonalization and the Lanczos algorithm [37]. These techniques calculate key physical quantities such as the localization length, density of states, and conductivity. Experimental validations of the Anderson model have been conducted using ultra-cold atom systems [16] and photonic lattices [53].
@@
Localization in disordered systems usually tends to be suppressed by an external magnetic field. For instance, weakly localized systems in two dimensions (2D) exhibit negative magnetoresistance due to a reduction of constructive interference of the electronic wave functions caused by an additional magnetic phase shift [13]. In the 2D Anderson model, a weak magnetic field reduces localization for states near the Landau band centers [8].
@@
A different class of systems, which are not randomly disordered but still lack perfect lattice periodicity, are quasi-periodic crystals. One of the most popular quasi-periodic 1D models is the Aubry-André (AA) model [10, 57, 21], which features periodically varying on-site energies with the period incommensurate with the lattice, meaning the ratio between the two periods is irrational. In the AA model, all states become localized when the potential strength exceeds a certain threshold. The model is closely related to the Harper model, which describes electrons in a two-dimensional periodic lattice under a uniform magnetic field; the Harper equation can be mapped onto the AA model [27]. The AA model can also be used to describe ultracold atoms in two incommensurate optical lattices. An experimental demonstration of the localization of bosonic atoms in such a system was reported in Ref. [49], offering direct evidence of the AA model predictions. Further experiments have investigated many-body localization in quasi-periodic systems [52].
@@
Several models related to the AA model have been proposed to explore more complex localization phenomena in quasi-periodic systems. One such model is the Soukoulis–Economou model [58], which introduces an additional potential term that breaks the self-duality of the original AA model. This modification allows the coexistence of both extended and localized states within the same system, separated by a mobility edge—critical energy value marking the transition between extended and localized behavior. Mobility edges have also been found theoretically in numerous other generalizations of the single-chain AA model [20, 15, 14, 43, 19, 63, 61]. Experimental evidence of a mobility edge in a 1D quasiperiodic optical lattice has been reported in Ref. [45].
@@
The Fibonacci chain [34, 1, 35, 30] provides yet another important example. In this 1D tight-binding model, the on-site potentials or hopping amplitudes follow a deterministic, quasi-periodic pattern based on the Fibonacci sequence. The resulting wavefunctions are critical, i.e. neither fully extended nor exponentially localized, instead decaying algebraically and exhibiting self-similar, fractal features. Interestingly, it has been shown that the AA model and the Fibonacci quasi-crystal are topologically equivalent, as one can be smoothly deformed into the other without closing any bulk energy gaps [38].
@@
In this paper, we construct a model of quasi-periodic crystals in which (i) incommensurability naturally arises from the geometry of the system and (ii) it is straightforward to study the effects of an external magnetic field on quasi-1D localization. Both goals can be achieved if one considers two coupled 1D chains of atoms, with the ratio $\rho$ of the lattice periods being an irrational number. We do not introduce any external on-site potentials, therefore, in contrast to the AA model and its multi-chain generalizations [26, 50, 42, 4, 33, 40], the incommensurability in our model affects only the hopping amplitudes.
Ladder models or “coupled moiré chains” with geometry similar to that considered here have been studied previously [62, 24]. Those works focused on Gaussian distance-dependent inter-chain hopping with short-range intra-chain coupling and demonstrated mobility edges in such systems. In contrast, we treat hoppings between sites in both chains on equal footing and consider exponential distance-dependent hopping for both intra- and inter-chain, with geometrically enforced periodic boundary conditions. We further relate the mobility edge to gaps in the energy spectrum, and study the effects of an external magnetic field on localization.
@@
For numerical analysis, we assume that the chains have different numbers of atoms but the same length, with periodic boundary conditions, and approximate $\rho$ by a ratio of two large coprime numbers. To characterize localization of the electronic wave functions, we use the the inverse participation ratio (IPR) [12, 60, 64]. The IPR takes values between zero and one, with the larger values corresponding to the more localized wave functions.
@@ II The Model
to closely approximate an irrational number. This can be achieved using the method of continued fractions [11, 56]. We represent the ratio (1) in the following form:
$\rho=a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{\ddots+\cfrac{1}{a_{r}}}},$ (2)
where $a_{0},a_{1},\ldots,a_{r}$ are integers [31]. In our analysis we will try to bring $\rho$ close to the golden ratio $\phi=(1+\sqrt{5})/2\simeq 1.618$, which corresponds to all $a_{i}=1$ ($i\geq 0$). Truncating this infinite sequence to $r$ terms leads to a rational approximation for $\phi$ called the $r$th convergent and denoted as $\phi_{r}$. The latter is given by a ratio of consecutive Fibonacci numbers:
$\phi_{r}=\frac{F_{r+1}}{F_{r}}.$ (3)
@@
see Fig. 1(b). Therefore, the hopping amplitudes take the following form:
@@
Note that $t_{A}$ and $t_{B}$ are invariant with respect to the lattice translations in the respective chains, i.e., they depend only on $|i_{A}-j_{A}|$ or $|i_{B}-j_{B}|$, but have different magnitudes if $N_{A}\neq N_{B}$. In contrast, the inter-chain hopping amplitudes are not translationally invariant.
Although all possible hopping terms are included in our model, in practice, for any given atom, only a few neighbors in both chains make a non-negligible contribution. The magnitudes of the inter-chain hoppings vary from atom to atom in a quasi-random fashion, producing a “geometrical disorder” in the system. We would like to mention here the study of the energy spectrum in a model of two coupled weakly incommensurate “Moiré chains” [18]. In this model, only the nearest-neighbor intra-chain hopping amplitudes (same in both chains) were considered and a different dependence of the inter-chain hoppings on the distance between the sites was assumed.
@@ II.1 Effects of magnetic field
We now include the effects of an external magnetic field in our tight-binding model, which can be achieved by using the Peierls substitution [28, 47]. In this method each hopping amplitude in Eqs. (5) and (6) is modified by a complex phase factor: $t_{\bm{r}\bm{r}^{\prime}}\xrightarrow{}t_{\bm{r}\bm{r}^{\prime}}e^{i\varphi_{\bm{r}\bm{r}^{\prime}}}$, where
@@ III Numerical Analysis
In periodic systems, Bloch’s theorem enables analytical calculation of the energy spectrum by assuming extended plane wave solutions for the wave function. However, our system does not repeat periodically; instead, it consists of $N_{A}+N_{B}$ atoms within a single “unit cell”, which coincides with the whole crystal. This non-periodicity means that extended plane-wave solutions are no longer valid and that Bloch’s theorem cannot be applied, resulting in $N_{A}+N_{B}$ distinct energy levels rather than energy bands. By increasing the size of the system and tuning the ratio $\rho$ to approximate an irrational number, we effectively render the system incommensurate. Since there is no analytical solution available for such a system, we must resort to numerical methods to find the wave functions and energy eigenvalues.
@@
In our model, following a considerable precedent in the previous studies of incommensurate 1D systems [10], we focus on using the golden ratio $\phi$ as the incommensurability parameter. To better represent real physical systems we select large values for $N_{A}$ and $N_{B}$. By choosing values of $N_{A}$ and $N_{B}$ in the thousands, we ensure a high degree of precision in approximating $\phi$ while maintaining practical relevance for numerical analysis. Namely, we use
@@
as determined from Eq. (3), i.e. set $N_{A}=F_{18}=2584$ and $N_{B}=F_{17}=1597$, where $F_{r}$ is the $r$th Fibonacci number. Selecting higher-order approximations does not qualitatively affect our results.
@@ IV Results
In our numerical calculations, we use $N_{A}=2584$ and $N_{B}=1597$ and diagonalize a $4181\times 4181$ Hamiltonian matrix for different values of the dimensionless lattice constant $a$ and the inter-chain separation $d$ (the lattice constant in the $B$ chain is given by $b=\rho a$). The effects of a magnetic field, which are described by the dimensionless parameters $\chi_{\parallel}$ and $\chi_{\perp}$, are considered in Sec. IV.4.
IV.1 Energy spectrum
@@
The evolution of the energy spectrum as a function of the inter-chain separation $d$ is shown in Fig. 5. At sufficiently large values of $d$ the spectrum is smooth and agrees with analytical results for decoupled chains. When $d$ is decreased, gaps begin to appear in the spectrum. Interestingly, these gaps seem to be located exactly at indices corresponding to the Fibonacci numbers $n=F_{r}$ or their simple combinations. This can be understood as an effect of “small denominators” in the perturbation expansions with respect to a weak inter-chain coupling, see Appendix A.
Figure 6: (Color online) The energy spectrum for different ratios $\rho=N_{A}/N_{B}$, chosen according to Eq. (20), for $N=4181$, $a=2$, and $d=1$. The vertical dashed line corresponds to $N_{A}=2584$ and $N_{B}=1597$.
Next, we examine how the incommensurability parameter $\rho$ defined in Eq. (1) influences the energy spectrum. We will focus on the case $N_{A}\geq N_{B}$. To explore the full range of $\rho$, we fix the total number of sites in the system, $N=N_{A}+N_{B}$, and allow $N_{A}$ and $N_{B}$ to vary under this constraint. In this parametrization, $\rho$ can be written as
$\rho=\frac{N_{A}}{\,N-N_{A}\,}$, (20)
with $N/2\leq N_{A}\leq N-1$ (equivalently, $1\leq N_{B}\leq N/2$), which restricts it to the interval $1\leq\rho\leq N-1$. Physically, varying $N_{A}$ in this way takes us from the “maximally commensurate” case where the two chains contain the same number of sites to the extreme limit where the $B$-chain contains only a single site. We plot the energy spectrum for each of these $(N_{A},N_{B})$ pairs in Fig. 6.
IV.2 IPR
@@
Our results for the coupled chains indicate that higher-energy states tend to be localized, exhibiting nonzero values of the IPR, whereas lower-energy states remain extended. Notably, the transition from extended to localized states is not continuous; instead, it appears to be abrupt and coincides with a gap in the energy spectrum. These findings suggest a strong correlation between the presence of energy gaps and the onset of localization in higher-energy states.
Plotting the IPR for individual eigenstates can be misleading, since the IPR of critical states also vanishes in the thermodynamic limit. To distinguish extended and localized states from critical ones, we therefore study how the IPR scales with the system size [22]. In general, the IPR scales as
$\mathrm{IPR}(N)\sim N^{-D},$ (21)
where $N$ is the total number of sites in the system, in our case $N=N_{A}+N_{B}$. Extended states are characterized by $D=1$, for which $\mathrm{IPR}\sim 1/N$, while localized states have $D=0$, corresponding to an IPR that remains finite as $N$ increases. Critical states exhibit intermediate scaling with $0<D<1$; although their IPR also vanishes in the thermodynamic limit, these states are neither fully extended nor localized.
We focus on the parameter regime with $a=2$, $d=1$, and $\rho$ approximating the golden ratio. In this case, the spectrum contains a high-energy region with nonzero IPR and a low-energy region with vanishing IPR. The two regions are separated by a mobility edge, as shown in Fig. 4(b). To characterize the nature of the states in each region, we compute the $\mathrm{IPR}$ averaged over a window of $0.01N$ eigenstates well inside each region, away from the mobility edge. We then analyze how the average $\mathrm{IPR}$ scales with the system size and extract the scaling exponent $D$, see Eq. (21). In accordance with Eq. (3), the system sizes are chosen such that $N_{A}$ and $N_{B}$ are consecutive Fibonacci numbers, ensuring that their ratio provides increasingly accurate rational approximations to the golden ratio.
Figure 10: Log-log plots of the average IPR vs the system size, fitted to a linear relation for $a=2$ and $d=1$. (a) Extended states in the low-energy region. (b) Localized states in the high-energy region.
In Fig. 10, we show the results of the IPR scaling analysis. In the low-energy region, we focus on the window of eigenstates centered at $n=N_{A}/2$ with width $0.01N$. From this analysis, we extracted a scaling exponent $D=0.99951\pm 0.00008$, which suggests that the states in this region are extended. In the high-energy region, we focus on the window of eigenstates centered at $n=N_{A}+N_{B}/2$ with width $0.01N$. In this case, we extracted the scaling exponent $D=-0.002\pm 0.008$, which is equal to zero within statistical uncertainty, suggesting that the states in this region are indeed localized. In both energy regions we find that scaling the system does not affect the nature of the eigenstates.
Figure 11: Log-log plots of the IPR vs the system size, fitted to a linear relation for $a=2$ and $d=1$. (a) Single state on the low-energy (extended) side of the mobility edge. (b) Single state on the high-energy (localized) side of the mobility edge.
@@
Next, we characterize the states at the mobility edge. The state with $n=N_{A}$ lies on the extended side of the mobility edge, while the state with $n=N_{A}+1$ lies on the localized side. In Fig. 11(a), we plot $\ln(\mathrm{IPR})$ of the state $n=N_{A}$ as a function of $\ln(N)$ and extract a scaling exponent $D=1.0\pm 0.1$, confirming that this state is indeed extended. In Fig. 11(b), we plot $\ln(\mathrm{IPR})$ of the state $n=N_{A}+1$ as a function of $\ln(N)$ and extract a scaling exponent $D=0.000\pm 0.003$, indicating that this state is indeed localized. Thus we see that the transition between extended and localized states at the mobility edge is abrupt, with no intermediate phase. This behaviour should be contrasted with the results of the AA model, where the wave functions at the critical point are neither fully localized nor extended, but exhibit irregular multifractal fluctuations across all length scales [54].
IV.3 Structure of the eigenstates
To illustrate the shape of the localized states, in Fig. 12 we plot the probability densities $|\psi_{n}(A,i_{A})|^{2}$ and $|\psi_{n}(B,i_{B})|^{2}$ using as an example the state with $n=2750$. The wave function exhibits sharply localized peaks in both chains in close proximity of each other. To characterize its spatial extent of the main peak, we approximate the probability density in each chain by an exponential function of the form
$|\psi(\theta)|^{2}=|\psi(\theta_{0})|^{2}\exp\left({-\frac{|\theta-\theta_{0}|}{\delta}}\right).$ (22)
Here, we used the angular coordinate $\theta$ to take advantage of the cylindrical symmetry of the system and introduced the angular localization length $\delta$. For convenience, we also define the linear localization length $\xi=R\delta$, allowing a direct comparison of the localization length with the lattice spacing.
To explore the detailed structure of the localized eigenstates, in Fig. 13 we plot the logarithm of the probability density over the angular coordinate $\theta$ for the $n=2750$ state. The wave function has two peaks, which nearly mirror each other in both chains. The shorter peaks correspond to a much smaller probability density, on the order of $e^{-10}$. We extract the localization lengths in the AA and $B$ chains by averaging the linear fits on each side of the main peaks, as shown in Fig. 13. For the $A$-chain, we find a localization length of $\xi_{A}/a=1.31\pm 0.02$, and for the $B$-chain we find $\xi_{B}/a=1.32\pm 0.01$. Thus, the localization lengths in the AA and $B$ chains are nearly identical.
(a) Refer to caption
(b) Refer to caption
Figure 12: (Color online) (a) “Bird’s-eye” view of the probability densities $|\psi_{n}(A,i_{A})|^{2}$ and $|\psi_{n}(B,i_{B})|^{2}$ in the localized state with $n=2750$, for $a=2$ and $d=1$. The smaller peaks have been scaled up in order to make them visible. (b) Logarithmic plot of the same probability densities without scaling.
Figure 13: (Color online) Logarithm of the probability densities in the AA and $B$ chains for the $n=2750$ state ($a=2$ and $d=1$). The data is fitted to a straight line on each side of the peak.
IV.4 Effects of magnetic field